Please solve it
x^2 +7x+12=0

Answer by nerdybill(7384)

x^2+7x+12=0
Factoring gives:
(x+4)(x+3) =0
.
Setting each term in parentheses to zero and solving gives you both solutions:
x+4 = 0
x = -4
.
x+3 = 0
x = -3
.
Therefore,
x = {-3, -4}